OpenStudy (anonymous):

A car travels at a constant velocity of 20.0 meters/second for 15 seconds. What is the power of the car if the initial force applied to it was 95 newtons?

OpenStudy (ybarrap):

We know that work is \(Force \times Distance\). We know that the Force is 95 N. But we need Distance to complete this problem. The car travels a distance of \(velocity \times time\) Meters because velocity is in meters/sec and time is in secs; so \( \dfrac{meters}{\cancel{sec}} \times \cancel{sec}\) = meters, which has the units of distance. The distance traveled in 15 seconds is : $$ 20~ \dfrac{meters}{\cancel{sec}} \times 15 \cancel{sec}=20*15~ meters $$ This tells us that the power of the car is \(Force \times Distance=95*20*15 ~Joules/second\)

OpenStudy (ybarrap):

Note that power has units of Energy per unit time. Energy is measured in Joules so Power is Joules per second.

OpenStudy (ybarrap):

Do you have any questions?